# GenTorrent: Scaling Large Language Model Serving with An Overlay Network

### v3

@@ Abstract
While significant progress has been made in research and development on open-source and cost-efficient large-language models (LLMs), serving scalability remains a critical challenge, particularly for small organizations and individuals seeking to deploy and test their LLM innovations. Inspired by peer-to-peer networks that leverage decentralized overlay nodes to increase throughput and availability, we propose GenTorrent, an LLM serving overlay that harnesses computing resources from decentralized contributors. We identify four key research problems inherent to enabling such a decentralized infrastructure: 1) overlay network organization; 2) LLM communication privacy; 3) overlay forwarding for resource efficiency; and 4) verification of serving quality. This work presents the first systematic study of these fundamental problems in the context of decentralized LLM serving. Evaluation results from a prototype implemented on a set of decentralized nodes demonstrate that GenTorrent achieves a latency reduction of over 50% compared to the baseline design without overlay forwarding. Furthermore, the security features introduce minimal overhead to serving latency and throughput. We believe this work pioneers a new direction for democratizing and scaling future AI serving capabilities.
@@ 1 Introduction
The rapid development and proliferation of open-source, cost-efficient large-language models (LLMs)–such as Google’s Gemma [2], DeepSeek-R1 [12], and Meta’s LLaMA series [43]–have significantly lowered the barrier for deploying AI systems. This progress makes it feasible to run advanced language models on personal workstations and local servers with consumer GPUs. Despite this accessibility, serving scalability remains a critical challenge for small organizations and individuals to deploy their LLM innovations. Individual researchers, university laboratories, and start-up companies often find it prohibitively expensive and technically challenging to deploy their models at scale, particularly when popularity surges quickly. For instance, some promising models initially provided excellent performance at low cost, but later suffered severe performance degradation when user demand suddenly increased [1]. Such challenges highlight the limitations of traditional cloud-based infrastructures, which are not always affordable or accessible for smaller organizations and individuals, and also pose significant privacy risks due to centralized user data management.
We propose GenTorrent, a decentralized overlay infrastructure designed for large-scale LLM deployment. Inspired by PlanetLab [11], a well-known peer-to-peer (P2P) network allowing researchers to deploy and test Internet applications across distributed nodes, GenTorrent is designed to harness computing resources from a diverse range of contributors, including individuals, academic institutions, and small organizations. Participants in GenTorrent volunteer their computing power and network resources, and in return gain the ability to deploy their own LLMs and AI-driven services on this shared, decentralized network.
GenTorrent provides the following main advantages:
@@
    Scalability, resiliency, and cost-efficiency. By distributing computational tasks across numerous decentralized nodes, GenTorrent can dynamically adapt to rapid fluctuations in demand, eliminating the performance bottlenecks commonly associated with centralized cloud deployments that depend on rigid, pre-defined service agreements. Idle computing resources (e.g., personal GPUs, university lab machines) can be utilized instead of expensive dedicated cloud instances.
@@
    Democratization of AI innovation development. GenTorrent significantly lowers the barrier to independent researchers, developers, and academic institutions being able to deploy and experiment with their LLM innovations at scale. Hence, it encourages innovation from a wider and more diverse contributor base.
@@
    Enhance user privacy. Conventional cloud deployments often centralize user prompts and associated private information, creating a single point of trust and potential vulnerability. GenTorrent addresses this by incorporating anonymous overlay networks, preventing model-serving nodes from easily linking user prompts to individual identities, thereby enhancing user privacy.
@@
    GenTorrent naturally inherits benefits that are common to decentralized overlay systems, such as service redundancy, geographic distribution of resources, and load balancing among heterogeneous servers.
GenTorrent thus represents a novel approach to deploying LLMs and other advanced AI services, democratizing large-scale AI capabilities and enhancing user privacy.
In designing and implementing GenTorrent, this work addresses four core research challenges that are critical for any decentralized LLM serving infrastructure:
@@
    Overlay network organization. We define the basic roles of overlay nodes in GenTorrent and detail the organization of the overlay network.
@@
    Anonymous communication. Decentralized model serving introduces privacy risks, as malicious overlay nodes might link user prompts to their identities. To mitigate this threat, we propose an anonymous communication service that protects user privacy by preventing the association of prompts with individual identities, even under overlay node churn.
@@
    Overlay forwarding for serving efficiency. GenTorrent enables model-serving nodes to forward user prompts across the network, allowing load balancing and key-value cache reuse. We design a decentralized forwarding mechanism that improves resource efficiency and achieves significant latency reduction while ensuring balanced workloads among model-serving nodes.
  4. 4.
    Decentralized verification of model serving. To maintain the quality and correctness of LLM serving, it is crucial to detect instances where model-serving nodes do not run their models as specified (e.g., deploying a smaller model). We develop decentralized verification to identify and mitigate such non-compliant behaviors.
To our knowledge, this is the first work to systematically study these core challenges for achieving LLM serving scalability via a decentralized overlay network independent of centralized cloud control. While further research and deployment of GenTorrent will likely reveal additional research challenges beyond the four discussed, we believe that these four issues are fundamental and critical to any decentralized LLM-serving infrastructure. We anticipate that this work will provide a foundation for future innovations and inspire continued research and development in decentralized network and system infrastructures for scalable AI serving.
We implemented GenTorrent’s core functions as a prototype consisting of decentralized nodes and tested the prototype in a public cloud without central control. This approach provided a controllable testbed, necessary as we do not have direct access to a large set of physical decentralized servers across the Internet. We let each node add synthetic latency to every packet for the wide-area Internet conditions. Upon acceptance of this paper, we will make the prototype implementation and data sets open-sourced. We plan to deploy and maintain GenTorrent at scale in the future.
This work focuses exclusively on the LLM serving problem in a decentralized overlay network and assumes model training is completed by model developers. We do not consider decentralized LLM training [48] or collaborative serving an LLM instance [5, 50, 33]. This work is also different from distributed LLM serving in a cloud with centralized scheduling [49, 41, 50, 33, 29].
The rest of this paper is organized as follows. Section 2 presents the system models and design objectives of GenTorrent. Section 3 details the design of GenTorrent. We provide the security analysis and evaluation in Section 4. The performance evaluation results are presented in Section 5. We discuss the related work in Section 6. We conclude this work and discuss possible future work in Section 7.
2 System Models and Objectives
Figure 1: Overview of GenTorrent.
This section presents the network and security models used by GenTorrent, whose overview is illustrated in Fig. 1. We also outline the design objectives of GenTorrent.
2.1 Network Model
GenTorrent operates as an overlay network composed of end hosts, each plays as one or more of the following roles:
  1. 1.
    User node represents end-users (clients) seeking access to LLM services while preserving privacy. To facilitate an anonymous overlay, each user node also agrees to forward packets for other users.
  2. 2.
    Model node executes LLM inference tasks to serve user requests. These nodes typically run on servers with GPU resources, contributed by participating organizations or individuals. Consistent with our system scope, we assume each model node serves one or more complete LLM instances. Collaborative inference across multiple nodes for a single request [5, 50] is not considered.
  3. 3.
@@ 2.1 Network Model
    Verification node. To maintain correct LLM serving and defend against security attacks, GenTorrent includes a committee of decentralized verification nodes, far fewer than users and model nodes. Each verification node runs on a server owned by an organization or individual with certain credentials, such as a well-recognized university research team. They are to ensure that each model node behaves legitimately and manage user and model node information, such as handling node registration and node reputation updates for model nodes. This committee runs a BFT consensus protocol [7] to ensure information correctness and consistency.
@@
A single physical device may operate in two or all roles in GenTorrent. All communications between nodes in GenTorrent are via TCP, secured with TLS. We assume every node is directly reachable via its IP address. User nodes may join or leave the network at any time and hence network churn can happen. Model and verification nodes are more stable but they still can fail or be compromised by attackers.
@@ 2.2 Threat Model and Assumptions
Consistent to many existing works [22, 20, 34], this work focuses on practical low-delay security and anonymity, rather than perfect security at all costs. We assume the following attacks by different nodes in GenTorrent, most of them are fundamental considerations in any decentralized model serving because they can be executed with small costs.
@@
  • A dishonest model node may drop user requests on purpose or reply with incorrect model results to mislead users. This includes manipulating outputs or substituting the agreed-upon model with a cheaper, lower-quality one. For example, the node claims to serve an 8B-parameter LLM but runs a 1.5B model instead to conserve resources. It will try to differentiate normal requests from users and verification requests to treat them differently. Even though a model node does not perform active attacks, it might attempt to link the received prompts and their user nodes. It might falsely claim resource exhaustion to refuse service.
@@ 2.3 Design Objectives
GenTorrent is designed to achieve the following objectives:
Performance: The key performance goals include providing low-latency LLM services with high scalability. In addition, GenTorrent aims to achieve load balancing, preventing individual nodes from being overwhelmed (i.e., receiving more requests than their capacity, leading to high completion times). Inspired by efficiency gains in centralized LLM serving [49, 41, 25], GenTorrent also targets key-value (KV) cache reuse across the overlay. This introduces a novel challenge specific to the decentralized context: efficient decentralized KV cache matching.
Security, privacy, and availability: GenTorrent is designed to ensure robust service availability despite network churn, node failures, and the adversarial behaviors identified in the threat model. Although malicious users may intentionally drop packets sent by others, clients can still achieve a high probability of successful request delivery due to path redundancy. Service integrity and security are maintained by the verification committee, which detects model nodes providing incorrect or low-quality responses and maintains accurate, consistent node reputations. GenTorrent aims to preserve user anonymity and protect the confidentiality of prompt content and model responses. It is designed to prevent model nodes or other users from linking user identities (e.g., IP addresses) to their requests.
3 System Design of GenTorrent
This section details the design of GenTorrent.
@@ 3.1 Overview
The overview of GenTorrent is illustrated in Fig. 1. The design of GenTorrent includes a number of protocols that allow nodes with the same or different roles to communicate with each other. We outline the protocols associated with each node type below.
@@
The user nodes form a dynamic overlay network. Upon joining GenTorrent, a user node registers its IP address and public key with the verification nodes, whose IP addresses and public keys are public information. It then downloads a list of active user and model nodes from a verification node, and the list is signed by more than 2/3 of the verification nodes [8]. For scalability in large deployments, verification nodes may divide the whole system into multiple regions and create a list of users and model nodes for each region. The requirement is that the number of users in the region should be sufficiently large to hide the requester’s identity, for example, >1000>1000 users, and the number of model nodes should be sufficiently large to achieve load balancing. GenTorrent allows a user to use other users as relays for anonymous communication: each user request will be forwarded along an overlay path to hide the address of the original sending user, and the LLM reply will also be forwarded along the anonymous path back to the requesting user.
@@
Each model node runs an instance of an LLM, which is deployed as a portable isolated environment, such as a container [32], provided by the LLM developer. One or more LLMs are deployed in the network, and each user request specifies which LLM it is requesting. A user request can be responded to by any model node that serves the same LLM as its request target. Model nodes use overlay forwarding to achieve load balancing and KV cache reuse among them.
@@
The committee of verification nodes uses a recent BFT consensus protocol [7] to ensure the information maintained by them is consistent, assuming <1/3<1/3 of the nodes are malicious. The committee maintains a list of user nodes registered to GenTorrent, including their public keys, and a list of model nodes in GenTorrent, including their public keys and reputation scores. The reputation score of a model node quantifies the quality of its LLM responses and will be defined in Section 3.4. Verification nodes periodically request the model nodes through the anonymous overlay and update their scores based on the responses.
3.2 Anonymous Overlay of Users
The anonymous overlay of users in GenTorrent allows each user node to request the LLM service without revealing its IP address. We summarize the requirements of the anonymous overlay as follows:
  1. 1.
    GenTorrent preserves user node anonymity, which requires that other nodes cannot link a prompt with the requesting user IP. However, anonymity of the model node is unnecessary: Users can link the IPs of the model node to the responses.
  2. 2.
    The confidentiality of prompts and responses should be preserved. The contents can only be visible to the requester and model node.
  3. 3.
    The anonymous overlay should not cause high overhead on user nodes. In particular, we do not want to use public-key cryptography for prompts and responses, which introduces a large computing overhead at relay nodes.
  4. 4.
    A single anonymous path is susceptible to relay failures under network churn and traffic analysis attacks. GenTorrent uses multiple paths to provide reliable message delivery.
We analyze typical overlay anonymous communication protocols [13, 22, 20, 34], including overlay Onion routing and sliced routing [22, 34], in the Appendix and find that none of them fit the above requirements.

### v4

# PlanetServe: A Decentralized, Scalable, and Privacy-Preserving Overlay for Democratizing Large Language Model Serving

@@ Abstract
While significant progress has been made in research and development on open-source and cost-efficient large-language models (LLMs), serving scalability remains a critical challenge, particularly for small organizations and individuals seeking to deploy and test their LLM innovations. Inspired by peer-to-peer networks that leverage decentralized overlay nodes to increase throughput and availability, we propose PlanetServe, an LLM serving overlay that harnesses computing resources from decentralized contributors. We identify four key research problems inherent to enabling such a decentralized infrastructure: 1) overlay network organization; 2) LLM communication privacy; 3) overlay forwarding for resource efficiency; and 4) verification of serving quality. This work presents the first systematic study of these fundamental problems in the context of decentralized LLM serving. Evaluation results from a prototype implemented on a set of decentralized nodes demonstrate that PlanetServe achieves a latency reduction of over 50% compared to the baseline design without overlay forwarding. Furthermore, the security features introduce minimal overhead to serving latency and throughput. We believe this work pioneers a new direction for democratizing and scaling future AI serving capabilities.
@@ 1 Introduction
The rapid development and proliferation of open-source, cost-efficient large-language models (LLMs)–such as Google’s Gemma [gemma], DeepSeek-R1 [deepseek], and Meta’s LLaMA series [touvron2023llama]–have significantly lowered the barrier for deploying AI systems. This progress makes it feasible to run advanced language models on personal workstations and local servers with consumer GPUs. Despite this accessibility, serving scalability remains a critical challenge for small organizations and individuals to deploy their LLM innovations. Individual researchers, university laboratories, and start-up companies often find it prohibitively expensive and technically challenging to deploy their models at scale, particularly when popularity surges quickly. For instance, some promising models initially provided excellent performance at low cost, but later suffered severe performance degradation when user demand suddenly increased [DeepseekSlow]. Traditional cloud-based infrastructures, which are not always affordable or accessible for smaller organizations and individuals.
We propose PlanetServe, a decentralized overlay infrastructure designed for large-scale LLM deployment. Inspired by PlanetLab [planetlab], a well-known peer-to-peer (P2P) network allowing researchers to deploy and test Internet applications across distributed nodes, PlanetServe is designed to harness computing resources from a diverse range of contributors, including individuals, academic institutions, and small organizations. Participants in PlanetServe volunteer their computing power and network resources, and in return gain the ability to deploy their own LLMs and AI-driven services on this shared, decentralized network.
PlanetServe provides the following main advantages:
@@
    Scalability, resiliency, and cost-efficiency. PlanetServe can dynamically adapt to rapid fluctuations in demand, eliminating the performance bottlenecks commonly associated with centralized cloud deployments that depend on rigid, pre-defined service agreements. Idle computing resources can be utilized instead of expensive dedicated cloud instances. PlanetServe increases scalability on both computation and networks. Hence, it can resolve the scalability problem for computation (e.g., large models), networks (e.g., small models), or a mixed situation.
@@
    Democratization of AI innovation development. PlanetServe significantly lowers the barrier to independent researchers, developers, and academic institutions being able to deploy and experiment with their LLM innovations at scale. Hence, it encourages innovation from a wider and more diverse contributor base.
@@
    Enhance user privacy. Conventional cloud deployments often centralize user prompts and associated private information, creating a single point of trust and potential vulnerability. PlanetServe addresses this by incorporating anonymous overlay networks, preventing model-serving nodes from easily linking user prompts to individual identities, thereby enhancing user privacy.
@@
    PlanetServe naturally inherits benefits that are common to decentralized overlay systems, such as service redundancy, geographic distribution of resources, and load balancing among heterogeneous servers.
PlanetServe thus represents a novel approach to deploying LLMs and other advanced AI services, democratizing large-scale AI capabilities and enhancing user privacy. Our position is that model serving today is mostly concentrated in large-scale cloud deployments, but the software ecosystem [mlc-llm] has already demonstrated that inference can be effectively deployed across a wide range of devices. This trend suggests that the future inference landscape should also be determined by cost efficiency, privacy, and availability. A decentralized infrastructure like PlanetServe is well-positioned to adapt to such a landscape: whether the dominant models are large or small, participants can contribute heterogeneous resources that align with their needs (e.g., low-latency private inference or specialized model serving).
In designing and implementing PlanetServe, this work addresses four core research challenges that are critical for any decentralized LLM serving infrastructure: 1) Overlay network organization. 2) Anonymous communication to protect user privacy. 3) Overlay forwarding for serving efficiency. 4) Decentralized verification of compliant model serving. To our knowledge, this is the first work to systematically study these core challenges for achieving LLM serving scalability via a decentralized overlay network independent of centralized cloud control.
We implemented PlanetServe’s core functions as a prototype consisting of decentralized nodes and tested the prototype in a public cloud without central control. This approach provided a controllable testbed, necessary as we do not have direct access to a large set of physical decentralized servers across the Internet. We let each node add synthetic latency to every packet for the wide-area Internet conditions. Upon acceptance of this paper, we will make the prototype implementation and datasets open-sourced. We plan to deploy and maintain PlanetServe at scale in the future.
This work focuses exclusively on the LLM serving problem in a decentralized overlay network and assumes model training is completed by model developers. We do not consider decentralized LLM training [ye2024openfedllm] or collaborative serving an LLM instance [borzunov2023petal, DistServe2024, SplitWise2024]. This work is also different from distributed LLM serving in a cloud with centralized scheduling [sglang24, Preble25, DistServe2024, SplitWise2024, CacheGen2024].
The rest of this paper is organized as follows. Section 2 presents the system models and design objectives of PlanetServe. Section 3 details the design of PlanetServe. We provide the security analysis and evaluation in Section 4. The performance evaluation results are presented in Section 5. We discuss the related work in Section 7. We conclude this work and discuss possible future work in Section 8.
2 System Models and Objectives
Figure 1: Overview of PlanetServe.
2.1 Network Model
PlanetServe operates as an overlay network composed of end hosts, each plays as one or more of the following roles:
@@ 2.1 Network Model
    User node represents end-users (clients) seeking access to LLM services while preserving privacy.
@@
    Model node executes LLM inference tasks to serve user requests. These nodes typically run on servers with GPU resources, contributed by participating organizations or individuals. Each model node serves one or more complete LLM instances. Collaborative inference across multiple nodes for a single request [borzunov2023petal, DistServe2024] is not considered.
@@
    Verification node. To maintain correct LLM serving and defend against security attacks, PlanetServe includes a committee of decentralized verification nodes, far fewer than users and model nodes. Each verification node runs on a server owned by an organization or individual with certain credentials, such as a well-recognized university research team. They are to ensure that each model node behaves legitimately and manage user and model node information. This committee runs a BFT consensus protocol [TendermintBFT] to ensure information correctness and consistency.
@@
A single physical device may operate in two or all roles in PlanetServe. All communications between nodes in PlanetServe are via TCP, secured with TLS. User nodes may join or leave the network at any time and hence network churn can happen. Model and verification nodes are more stable but they still can fail or be compromised by attackers.
2.2 Cost Advantage and Incentive Model
Cost advantage of PlanetServe. The cost advantage of PlanetServe compared to commercial public cloud is mainly from two facts. First, computation resources remain dramatically underutilized in organizations. Hence, the cost of contributing servers to PlanetServe is very low. Many organizations deployed GPU servers for ML training, but the servers sit idle outside of training times. It has been reported that the average server utilization rate is between 12%-18% of capacity [underulti]. 10 million servers are estimated to sit completely idle, representing \$30 billion in wasted capital [underulti]. These idle servers can be contributed to PlanetServe and form a large-scale serving infrastructure. Second, using community-contributed GPUs for inference can dramatically cut costs compared to commercial cloud platforms. Measurements show that the energy cost of running LLM inference on GPUs such as the V100 and A100 is relatively modest [samsi2023words], implying that the high hourly rates of public cloud instances are driven largely by infrastructure and pricing overheads rather than intrinsic hardware or energy expenses. Recent work on scalable inference similarly reports $20\times$ lower cost when serving LLMs on consumer GPUs versus AWS instances [ScaleLLM]. In addition, many consumer GPUs or GPUs that have been given away from training can still serve moderately sized LLMs (7B–13B parameters).
In summary, prior measurements and cost analyses suggest that LLM serving and testing using private hardware can be an order of magnitude cheaper than compared to public cloud if the hardware is fully utilized. However, scaling private hardware is challenging, and hardware utilization is very low. PlanetServe addresses both the scalability and under-utilization problem of private hardware by allowing different organizations to contribute their resources and share them with each other.
PlanetServe uses a reputation-based incentive model to encourage participation and cooperation. The incentive model only applies to the organizations that contribute model nodes to the system. There is no system-wide incentive mechanism for users. However, model deployers can use their own policies to regulate users who can access their LLMs.
Incentive for participation. Each organization participates in the system by contributing a number of model nodes. Model nodes from the same organization share the same reputation score $\lambda$, which will be updated during the system execution, as detailed in Sec. 3.4. If the reputation score is above a threshold, the organizer is allowed to deploy their own LLM to the system. To measure the resources it is allowed to use for deploying its LLM, each organization also has a contribution credit, which reflects the resource time it contributes to the system. The contribution credit is calculated proportionally to the cost of renting servers from a public cloud. For example, if an organization has contributed 5 servers that have been serving for 30 days in PlanetServe, it can deploy its LLM to PlanetServe that runs on 30 servers with similar computing resources for 5 days. All reputation scores and contribution credits are assigned and updated by the committee of verification nodes running a consensus protocol.
Incentive for cooperation. If model nodes do not follow the PlanetServe protocols in model serving and request forwarding, their reputation scores will be reduced, which will hurt their eligibility to use other resources in PlanetServe. The verification node committee is responsible for ensuring that the model nodes correctly run the models they promise to serve and for maintaining the reputation scores of the model nodes, as detailed in Sec. 3.4.
@@ 2.3 Threat Model and Assumptions
Consistent to many existing works [IS-NSDI07, LAP12, garlic], this work focuses on practical low-delay security and anonymity, rather than perfect security at all costs. We assume the following attacks by different nodes in PlanetServe.
@@
  • A dishonest model node may drop user requests on purpose or reply with incorrect model results to mislead users. This includes manipulating outputs or substituting the agreed-upon model with a cheaper, lower-quality one. For example, the node claims to serve an 8B-parameter LLM but runs a 1.5B model instead to conserve resources. It will try to differentiate normal requests from users and verification requests to treat them differently. Even though a model node does not perform active attacks, it might attempt to link the received prompts and their user nodes. It might falsely claim resource exhaustion to refuse service. A malicious model node may also be curious about sensitive prompt content, such as users’ personal data or proprietary information.
@@ 2.4 Design Objectives
PlanetServe is designed to achieve the following objectives:
Performance: The key performance goals include providing low-latency LLM services with high scalability. In addition, PlanetServe aims to achieve load balancing, preventing individual nodes from being overwhelmed.
Security, privacy, and availability: PlanetServe aims to preserve user anonymity and protect the confidentiality of prompt content and model responses. It is designed to prevent model nodes or other users from linking user identities (e.g., IP addresses or public keys) to their requests. For privacy-sensitive workloads, PlanetServe leverages NVIDIA Confidential Computing on Hopper and Blackwell GPUs to run inference inside trusted execution environments (TEEs), ensuring that model nodes cannot access prompts or proprietary information. [NVH100, nvidia-secure-ai-blackwell-hopper]. In addition, PlanetServe should provide high service availability under node churn and attacks and ensure model node provide accurate responses.
3 System Design of PlanetServe
This section details the design of PlanetServe.
@@ 3.1 Overview
The overview of PlanetServe is illustrated in Fig. 1.
@@
The user nodes form a dynamic overlay network. Upon joining PlanetServe, a user node registers its current IP address and public key with the verification nodes, whose IP addresses and public keys are public information. The public key serves as the identifier. It then downloads a list of active user and model nodes from a verification node, and the list is signed by more than 2/3 of the verification nodes [PBFT99]. Verification nodes may choose to divide the whole system into multiple regions and create a list of users and model nodes for each region, only when the number of users in each region is sufficiently large to hide the requester’s identity, for example, >1000>1000 users [Intruder, hordes00]. PlanetServe allows a user to use other users as relays for anonymous communication: each user request will be forwarded along an overlay path to hide the address of the original sending user, and the LLM reply will also be forwarded along the anonymous path back to the requesting user.
@@
Each model node runs an instance of an LLM, which is deployed as a portable isolated environment, such as a container [Container2017], provided by the LLM developer. One or more LLMs are deployed in the network, and each user request specifies which LLM it is requesting. A user request can be responded to by any model node that serves the same LLM as its request target. Model nodes use overlay forwarding to achieve load balancing and KV cache reuse among them.
@@
The committee of verification nodes uses a recent BFT consensus protocol [TendermintBFT] to ensure consistency. The committee maintains a list of user nodes registered to PlanetServe, including their public keys, and a list of model nodes in PlanetServe, including their public keys and reputation scores. The reputation score of a model node quantifies the quality of its LLM responses and will be defined in Section 3.4. Verification nodes periodically request the model nodes anonymously and update their scores based on the responses.
3.2 Anonymous Overlay and Content Privacy
The anonymous overlay of users in PlanetServe allows each user node to request the LLM service without revealing its IP address. We summarize the requirements of the anonymous overlay as follows: 1) User node anonymity: other nodes cannot link a prompt with the requesting user identity. 2) Confidentiality of prompts and responses: the contents can only be visible to the requester and model node. 3) Low overhead on user nodes. In particular, we do not want to use public-key cryptography for prompts and responses, which introduces large overhead at relay nodes. 4) A single anonymous path is susceptible to relay failures under network churn. PlanetServe uses multiple paths to provide reliable message delivery.
We analyze typical overlay anonymous communication protocols [Tor04, IS-NSDI07, LAP12, garlic], including Onion routing and sliced routing [IS-NSDI07, garlic], in the Appendix and find that none of them fit the above requirements. If we apply Onion to the overlay network in PlanetServe, there are a few main challenges. 1) The first layer of Onion nodes (called guards in Tor) will definitely know the sending user of a request, which should be avoided in PlanetServe. 2) The failure probability of an Onion path increases exponentially with the path length, especially for an overlay network of users with a high churn rate. 3) Each relay needs to perform public-key decryption on the entire message, introducing non-trivial overhead for users.